OpenStudy (anonymous):

9. Which sum or difference is equivalent to the following expression? (1 Point)

OpenStudy (anonymous):

\[\frac{ 2x+3 }{ 4 }\]

OpenStudy (anonymous):

is there choices?

OpenStudy (anonymous):

Yes hold on for a sec

OpenStudy (anonymous):

A.\[\frac{ x }{ 2 }\]+\[\frac{ 3 }{ 4 }\] B.\[\frac{ x }{ 2 }-\frac{ 3 }{ 4 }\] C.\[\frac{ 3x }{ 4 }+\frac{ 3 }{ 4 }\] D.8x+24

OpenStudy (anonymous):

if you plug in a random # say 3 the original would = 2.25 and the only one that would be the same is a

OpenStudy (anonymous):

I was thinking a, but thanks alot .

OpenStudy (anonymous):

yep no problem